Find the y-intercept for the equation 2x+4y=20
Answer by bucky(2189)

You can solve this quite easily if you think about it. Ask yourself what the x value is
for any point on the y-axis. It may become clear that any point on the y-axis will have a
corresponding value of zero for x.
.
So in this equation all you need to do is to set the value of x equal to zero and solve
for the corresponding value of y.
.
The equation you were given is:
.
2x + 4y = 20
.
When you set x equal to zero, the term 2x becomes zero and therefore, it drops out. The
equation is then:
.
4y = 20
.
Solve for y by dividing both sides by 4 and you get:
.
y = 20/4 = +5
.
Therefore, the y intercept is the point (0, +5) and it is on the y-axis at a value of +5.
.
Hope this clarifies the problem for you and helps you to see how to do the problem.